### v1

# 2D Entanglement from a Single Motzkin Spaghetto

Abstract
While highly entangled ground states of gapless local Hamiltonians have been known to exist in one dimension, their 2D counterparts were only recently found, with rather sophisticated interactions involving at least 4 neighboring degrees of freedom. In this letter, a 2D lattice system interacting only between two nearest neighboring sites is proposed, which has a unique ground state with volume scaling of entanglement entropy. The system undergoes an entanglement phase transition with enhanced entanglement entropy at the critical point compared to the previous construction from coupling orthogonal arrays of one-dimensional chains, but the spectral gap in the highly entangled phase decays faster. A tensor network representation is composed of tensors with lower rank, and has a similar global geometry.
Introduction—Entanglement entropy (EE) is a measure of how entangled a quantum state is, or how different it is from a classical state. In that sense, the ground state (GS) of a locally interacting quantum many-body system separated by a spectral gap from the first excited state is not so different from a classical state, at least compared to generic eigenstates in the Hilbert space [1], as the EE of a subsystem grows with the size of the boundary instead of the bulk. This is known as the area law of EE [2], and is expected to hold in any spatial dimension. It is intuitive because massive quasiparticles propagate at a finite speed, resulting in exponentially decaying correlations. So only degrees of freedom within a finite range are effectively entangled across a boundary and contribute to the bipartite EE. The idea has been made rigorous into a proof for 1D systems using Lieb-Robinson bounds [3]. However, in 2D progress towards establishing a theorem has only been made for certain class of Hamiltonians known as frustration free [4], to which the model proposed in this letter belongs.
What the area law does not tell us is how entangled a GS can be when the spectral gap closes, as happens in quantum criticality. There, the EE typically scales with an extra factor logarithmic in the subsystem size beyond the area law. This has been demonstrated with conformal field theories for (1+1)D critical points [5, 6], and with the Widom conjecture for gapless free fermions with a Fermi sea in arbitrary dimensions [7]. That was perhaps not so surprising as correlation is known to have a power law decay at critical points. But what was a little striking was that when the spectral gap decays even faster with the system size, for example at an exponential rate as opposed to the polynomial decay for critical systems, the EE can actually have a volume scaling. This was first exhibited in an inhomogeneously interacting XX spin chain [8, 9], where the spins can be approximated to form Bell pairs symmetrically across the center in the GS, leading to maximal EE. The model was hence referred to as the rainbow chain.
Similar examples of highly entangle GSs with more exact descriptions were made possible with frustration free Hamiltonians [10], that are also translationally invariant (broken only by the finite boundary). For those Hamiltonians, the GS is simultaneously the lowest energy eigenstate of each local interaction operator, such that a lot can be known analytically about the GS by diagonalizing a few-body Hamiltonian. By mapping spin-1 configurations to discrete random walks, the GS of the critical Hamiltonian in Ref. [10] can be given as a superposition of Motzkin paths, those that start and end at hight zero, without ever passing below it in the middle. Using combinatorial results, Bravyi et al. showed that the GS of the Motzkin chain has the same scaling behavior for EE, correlation and spectral gap as typical critical systems. Movassagh and Shor enlarged the local Hilbert space of the Motzkin chain with a color degree of freedom, which has a power law scaling of EE that outgrows the spin sector [11]. These constructions were later generalized to spin-$\frac{1}{2}$ systems, with smaller local degrees of freedom, but three-body interaction involving next-nearest neighboring sites [12, 13], giving its name Fredkin chain. To actually realize volume scaling in the GS, these Hamiltonians have to be deformed in a frustration free way, such that the GS superposition favors higher paths which allow more Bell pairs across the bipartition boundary [14, 15, 16]. As the deformation parameter is tuned, the GS undergoes an entanglement phase transition from area law to volume scaling.
It is not so straightforward to generalize the physical picture of Bell pair from 1D to 2D, as entanglement in two different directions turns out to compete against each other. A naive generalization to the rainbow chain was made in Ref. [17], only to end up with a lattice of Hausdorff dimension one embedded in 2D space. The trade-off between entanglement in orthogonal directions was also exemplified in colored generalizations of the toric code model [18, 19], where enhanced entanglement along the bipartition boundary results in anomalous topological EE of the GS. Another obstacle is a well-defined notion of height in 2D, when height differences can be counted along different paths connecting two points in the lattice. The prescription to fix both difficulties is to couple two orthogonal arrays of 1D Motzkin or Fredkin chains with appropriate vertex or tiling rules, such that a Coulomb gauge emerges to define an unambiguous height function [20, 21]. Moreover, the coupling leaves the Hamiltonian frustration free, and do not interfere with the interactions along each entangled chain. So unlike the anisotropic coupled wire construction of topological superconductors [22] where the inter-wire coupling opens up a gap, the system remains gapless. The resulting GSs are superpositions of discrete random membranes with a hard wall constraint at zero height, weighted by the deformation parameter to the power of the volume under the surfaces. The same entanglement phase transition occurs, except that the reduced height fluctuation in higher dimensions results in a logarithmic violation to the area law at the critical point instead of power law as in the 1D models.
While the six- or 19-vertex and lozenge tiling realizations appear to be the natural generalization of the highly entangled Motzkin (resp. Fredkin) chain to 2D, the support of the local Hamiltonian operators grows from two (resp. three) in 1D to four (resp. six). So while the paradigm in principle works for higher dimensions, the interactions become increasingly less realistic. An alternative loop-soup construction was proposed in Ref. [18], where the GS is a superposition of random loop configurations on a 2D lattice, and on each loop lives a periodic Motzkin chain [23]. However, due to the randomness of the loop configurations, it is not clear how the EE for any bipartition would actually scale with the subsystem size. Moreover, a naive generalization of the Motzkin Hamiltonian to periodic boundary conditions leaves the GS with exponential degeneracy, making a combination with exotic EE scaling trivially exist. To avoid both of those complications, we focus on a 2D lattice filled by winding one open 1D chain. The interaction is turned on for bonds along the chain, so it remains two- or three-body. Since the location of the entangling chain is fixed, we are able to argue from the 1D results that the EE scaling is asymptotically the same as the coupled-chain construction, except at the critical point, and the entanglement is isotropic when the subsystems are partitioned in different directions. Furthermore, a tensor network (TN) representation of the new GS using the previously obtained TNs for the 1D chains [24, 25] has the same global geometry as the TN for the GS of the six-vertex coupled Fredkin chains [26].
Motzkin chains—The $q$-deformed bicolor Motzkin chain [14] has a five-dimensional local Hilbert space, with spin up and down, both in either red (r) or blue (b) color, and a colorless spin zero. The Hamiltonian is expressed in terms of projectors $P_{j}^{c}=|L^{c}\rangle_{j,j+1}\langle L^{c}|_{j,j+1}+|R^{c}\rangle_{j,j+1}\langle R^{c}|_{j,j+1}+|F^{c}\rangle_{j,j+1}\langle F^{c}|_{j,j+1}$, where
$\begin{split}|L^{c}\rangle_{j,j+1}=&\frac{1}{\sqrt{1+q^{2}}}\left(|\uparrow^{c}_{j}0_{j+1}\rangle-q|0_{j}\uparrow^{c}_{j+1}\rangle\right),\\ |R^{c}\rangle_{j,j+1}=&\frac{1}{\sqrt{1+q^{2}}}\left(|0_{j}\downarrow^{c}_{j+1}\rangle-q|\downarrow^{c}_{j}0_{j+1}\rangle\right),\\ |F^{c}\rangle_{j,j+1}=&\frac{1}{\sqrt{1+q^{2}}}\left(|\uparrow^{c}_{j}\downarrow^{c}_{j+1}\rangle-q|0_{j}0_{j+1}\rangle\right).\end{split}$ (1)
Each of the projection operator has a local lowest energy eigenstate corresponding to the direction orthogonal to the above three vectors. So the GS of a Hamiltonian consisting of such projectors acting on all neighboring pairs of sites in the chain is a weighted superposition of random walks if the up, down and 0 spins are respectively mapped to up, down and flat moves. In addition to the color matching encoded in $|F^{c}\rangle_{j,j+1}$ between up and down moves when they are pair created from spin zeros, an energy penalty needs to be introduced by the projector $C_{j}^{c_{1},c_{2}}=|\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}\rangle\langle\uparrow^{c_{1}}_{j}\downarrow^{c_{2}}_{j+1}|$ for $c_{1}\neq c_{2}$ in order to enforce color matching between nearest up-down pairs in the GS configurations. The remaining GS degeneracy can be lifted by fixing the boundary conditions with the Hamiltonian $H_{\partial}=\sum_{c=\mathrm{r,b}}|\downarrow^{c_{1}}\rangle\langle\downarrow^{c_{1}}|_{1}+|\uparrow^{c_{1}}\rangle\langle\uparrow^{c_{1}}|_{2N}$. In summary, the Motzkin Hamiltonian is given by
$H=H_{\partial}+\sum_{j=1}^{2N-1}\left(\sum_{c=\mathrm{r,b}}P_{j}^{c}+\sum_{c_{1}\neq c_{2}}C_{j}^{c_{1},c_{2}}\right).$ (2)
It has a unique GS as a superposition of spin configurations where there is always at least as many up spins as down spins to the left at any point along the chain, and the colors of the nearest up-down spins pairs are either both red or both blue, in a nested fashion. If up (resp. down) spins are mapped to $45^{\circ}$ upward (resp. downward) moves with unit horizontal lengths, and the area enclosed between a Motzkin path $m$ and the horizontal axis at height zero is denoted $A(m)$, then up to an overall normalization factor, the GS can be expressed as
$|\mathrm{GS}\rangle=\sum_{m\in M_{2N}}q^{A(m)}|m\rangle,$ (3)
where $M_{2N}$ denotes the set of bicolor Motzkin paths of length $2N$ with color matching.
2D lattice—Since the 2D model is constructed by winding the 1D chain, we only have to specify the location of the 1D chain on a 2D grid. This is shown in Fig. 1. The spin degrees of freedom reside only on the edges of the square lattice covered by the winding 1D chain, and interactions are only turned on between adjacent spins on the chain. Depending on the location of the spin in the square lattice, it can be either between east-west (resp. north-south) neighbors for the upper and lower (resp. left and right) quadrants, or neighbors around a corner near the two diagonals.
Figure 1: A twisted Motzkin chain of length $2N=L(L+2)$ occupying an $L\times L$ square lattice, shown here for $L=8$. The spin degrees of freedoms live on the edges of the square lattice, as are labeled in the figure. Four symmetric bipartition cuts are shown in different colors.
To study the isotropic EE, four symmetric bipartition cuts are shown in Fig. 1. They separate the spins into two subsystems consisting of increasingly larger segments of size $4n-a$, for $n=1,2,\cdots,L/2$, with $a=0,1,2,3$ corresponding respectively to the yellow, blue, green and red cut. Starting from the center, segments within one subsystem in the 2D lattice on a given side of the cut alternate between the first and second halves of the 1D chain when it is stretched straight. Obviously the scaling behavior of bipartite EE for any of the cuts will not differ qualitatively, so from now on we focus on the cut along the main diagonal with $a=0$, which results in a bipartition shown in Fig. 2.
Figure 2: The two subsystems resulting from the yellow cut in Fig. 1. The northeast subsystem is represented with the thickened black lines, whereas the southwest is marked by thin gray lines. The configuration depicted corresponds to the probabilistically dominant Motzkin path in the GS superposition for the $q>1$ phase, to illustrate the entanglement across subsystems. Typical height configuration for at the critical point is depicted in blue.
Scaling of EE—As analyzed in Ref. [14], the EE decomposes into a leading contribution from the color degree of freedom, and a subleading contribution from the spin sector. Only up-down spin pairs belonging to different subsystems account for the former. In the Motzkin path picture, this means that only the net height change of a subsystem, or the surplus of up spins matters. For the 1D system studied there, the leading contribution to the EE is therefore proportional to the height at the center of the Motzkin path, giving the three distinct scaling behaviors in the two different phases and the critical point.
Here, the situation is a little more complicated, as segments belonging to two subsystems alternate along the chain. But we can get lower or upper bounds for the two phases with the following argument. For $q>1$, the net height change within each segment scales linearly with its length, at least for segments long enough which are the overwhelming majority in the asymptotic limit. If there is a dip in the middle of the segment, it would only increase the entanglement between the two subsystems, as the immediate neighboring segment on the left belongs to the other subsystem. For down spin to actually form up-down pairs with up spins from the next neighboring segment to the left, coming from the same subsystem, the probability would be of the next subleading order. Therefore, assuming the height to monotonically increase for the first half of the chain gives a lower bound on the EE, which is already growing linearly following the proof for the 1D case [14].
For $q<1$, the net height fluctuation stays at order $\mathcal{O}(1)$ anywhere along the chain. In the best case scenario for accumulating EE, the height changes are positive for all the segments in one subsystem, and negative in the other. Since there are in total $L/2$ segments of each, the EE scales at most with the linear size of the lattice, which means the area law.
The $q=1$ case calls for more careful analysis, where the hard wall constraint from below determines the shape of typical Motzkin paths. Recall that the most probable height of a Motzkin path that starts from zero height scales as $h\approx\sqrt{x/3}$ after $x$ steps [10]. Following a completely analogous enumeration of discrete paths using the reflection method, and then taking the asymptotic limit, it can be shown that a Motzkin path that starts at height $h_{0}$ typically ends at height $h_{1}\approx\sqrt{(x-h_{0}^{2})/3}$ after $x$ steps. Iteratively applying this formula shows the height change within each segment scales as $\sqrt{L}$. Since there are order $L$ sufficiently long segments, the total EE scales as $L^{3/2}$, different from the $L\log L$ scaling of the couple chain constructions [20, 21].
The subleading contribution to the EE is from the fluctuation in the height of Motzkin paths, which is also the sole source of entanglement for a monochromatic version of the model. A Schmidt decomposition of the reduced density matrix shows that the two subsystems only depend on each other via the total spin within one subsystem, such that the total spin of the whole chain is zero. Moving from the 1D system to our 2D construction, nothing changes except the subsystem is distributed into disconnected segments. So dimensionality of the reduced density matrix is still the number of possible values of the total spin. However, in the $q>1$ (resp. $q<1$) phase, the probability of the total spin, or height change decays exponentially with the deviation of the area underneath a Motzkin path from the most probable one that is of highest height (resp. completely flat) [14, 15]. So only paths with a height change at finite distance away from the typical one effectively matters, given a bounded EE independent of system size for both phases. When $q=1$, there is more fluctuation. Although entropic repulsion pushes the average height to order $\sqrt{N}$ above the hard wall at zero height, the variance of the Gaussian free field remains to scale linearly with the total length of the segments, making the EE scale as $\log N$ [10]. Note that in a monochromatic version of the 2D model, all three scalings are sub-area law, which is unique to our construction of a 2D lattice by winding a 1D chain. The EE scalings for the two different 2D models are compared in Table 1 for different phases.
Table 1: Comparison of the EE scalings with linear system size $L$ between the two different 2D constructions.
Model Phase Coupled chains Twisted chain
$q>1$ $L^{2}$ [20, 21] $L^{2}$
Colored $q=1$ $L\log L$ [20, 21] $L^{3/2}$
$q<1$ $L$ [20, 21] $L$
$q>1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
Colorless $q=1$ $L$  [^1] This is an upper-bound from the zeroth order Rényi entropy. The random membrane when bipartitioned gives a Motzkin path on the boundary, giving the dimensionality of the reduced density matrix. The same upper-bound was also obtained for six-vertex models with domain wall boundary conditions using the exact enumeration of alternating sign matrices [27]. $\log L$
$q<1$ $\mathcal{O}(1)$ $\mathcal{O}(1)$
Spectral gap—Unlike the scaling of EE discussed above, which requires somewhat different analyses than the 1D system due to the more complicated bipartition, upper-bounds on the scaling of the spectral gap can be straightforwardly deduced from the 1D scaling simply by substituting the linear system size with $N=L(L+2)$. The results are summarized in Table 2. This should be compared with the EE scaling in Table 1. In particular, the $q>1$ phase of the colored model shows that the same EE scaling does not imply the same scaling of spectral gap, and vice versa as shown by the critical point of the colored model and the $q>1$ phase of the colorless one.
Table 2: Comparison of upper-bounds on the scaling of spectral gap with the linear system size $L$.
Model Phase Coupled chains Twisted chain
$q>1$ $e^{-\alpha L^{3}}$ [20] $e^{-\beta L^{4}}$ [28, 16]
Colored $q=1$ $L^{-\gamma}$  [^1] Conjectured in Ref. [26]. $L^{-\delta}$ [11]
$q<1$ $\mathcal{O}(1)$ 11footnotemark: 1 $\mathcal{O}(1)$ [29]
$q>1$ $e^{-\epsilon L^{2}}$ [20] $e^{-\zeta L^{2}}$ [28, 16]
Colorless $q=1$ $L^{-\eta}$ 11footnotemark: 1 $L^{-\theta}$ [10]
$q<1$ $\mathcal{O}(1)$ 11footnotemark: 1 $\mathcal{O}(1)$  [29]
TN—The GS of the 1D Motzkin chain has a holographic TN representation [24], where each physical spin maps to a column of tensors sequentially contracted with neighbors above and below. The depth of the columns grows linearly from the boundaries of the chain to the center. Twisting the TN from the center results in a 2D TN for the GS of the present model, which has a pyramid shape. The geometry of the TN is globally the same as the 3D TN for the GS of the coupled Fredkin chain model [26]. But the tensors in the bulk have only four internal legs instead of six. The number of bonds crossing the bipartition cut gives the correct scaling of the EE for the $q>1$ phase of the colored model. At the critical point of the colorless model, the TN can be renormalized such that the depth of the columns grow logarithmically instead, giving the correct EE scaling in that case [25].
Conclusion—IN this letter, a frustration free 2D lattice Hamiltonian with exactly solvable GS that has volume scaling of EE is constructed by twisting the highly entangled Motzkin chain. The interaction is significantly simpler than the previous and only model of such highly entangled 2D systems realized by coupling arrays of highly entangled chains. Therefore it has relatively better prospect in experimental realization and applications in quantum control and simulation. A spin-$1/2$ version of the model can be constructed by twisting a Fredkin completely analogously. At the moment, it is not clear if similar construction works for 3D systems, whereas the coupled chains construction in principle works in any dimension via tiling. These alternative examples of area-law violating GSs provide a better understanding of the interplay between EE and spectral gap scaling, or the tradeoff between quantum computing resource and robustness against noise. The ultimate goal in this direction is to generalize the area law to gapless ground states to get a relation between the rate of decay in spectral gap and growth of EE with the system size.
Acknowledgements.
I thank Shankar Balasubramanian, Soonwan Choi, Israel Klich, Ethan Lake and Olai Mykland for discussions.